# LLMs Capture Urban Science but Oversimplify Complexity

### v2

@@ 3 Discussion
This study reveals a fundamental duality in the capabilities of current Large Language Models within urban science. They excel at reproducing the abstract, mathematical forms of established theories but falter in capturing the statistical complexity and fidelity of real world urban systems. The systematic evaluation across different scales demonstrates that while LLMs can generate data that is structurally ”form similar” to urban theories, this apparent competence masks significant deficiencies in diversity and parametric accuracy. These findings position LLMs not as high fidelity physical simulators, but as powerful engines for exploring the socio-cultural patterns embedded in human language, a role that offers both profound opportunities and critical limitations for scientific inquiry.
@@
The initial finding that LLMs consistently generate ”mirage cities” with low diversity and systematically distorted parameters (e.g., underestimated scaling exponents) points to a core limitation. This phenomenon, where models produce archetypal representations of cities, a tendency sometimes described as ”group flattening” [31], aligns with observations of geographical and social bias in other domains [32, 33]. This interpretation suggests that these distortions are not random errors but are symptomatic of the LLMs’ nature as statistical pattern matchers rather than causal world models. Lacking a deep, generative understanding of the mechanisms that produce urban phenomena, they instead reproduce the most dominant statistical patterns from their training data, smoothing over the granular heterogeneity and outliers that are often crucial drivers of real world dynamics. This tendency, also noted in other applications of LLMs for generating synthetic populations [31], poses a significant risk for empirical research that relies on capturing the full spectrum of urban complexity.
@@
The investigation into data fidelity reveals that LLMs are not static instruments but are malleable tools sensitive to user interaction and model specific characteristics. Our finding that prompt engineering can partially mitigate deviations by steering the model toward a specific context underscores the importance of the human AI interface in scientific applications [30, 31]. The quest is not for a single, perfect prompt, but for a methodology of iterative refinement where the prompt itself is a key parameter in the experimental design [42]. Furthermore, the observed performance differences across the five evaluated LLMs highlight that model architecture and training strategies are critical variables. This suggests that future improvements in data fidelity will depend on both more sophisticated interaction protocols and the rapid advancements in the models themselves [20].
The limitations of LLMs also unveil a novel scientific opportunity. Their tendency to generate theoretically ”clean” data, stripped of real world noise, makes them powerful instruments for theoretical exploration [29, 34, 35, 31, 33]. Urban scientists can now generate vast, fine grained datasets under specific theoretical assumptions, creating virtual laboratories to explore a theory’s logical consequences or test its sensitivity to certain parameters. This moves beyond simply validating models against existing data and toward a new mode of inquiry focused on understanding the internal consistency and emergent properties of theories themselves. This approach directly addresses the challenge of how AI can generate new knowledge for urban studies, by providing a scalable method to rigorously interrogate theoretical constructs in a controlled environment.
@@
A distinct capability of LLMs is revealed when they are tasked with navigating less formalized, qualitative theories like urban vitality. Here, the LLM is not solving a physical equation but is prompted to simulate a ”social consensus” by synthesizing the vast amount of information about cities embedded in its training data, from academic texts to public discourse. The result is not a ”discovery” of a new law, but the generation of a plausible, structured set of relationships that can provide urban scientists with novel, testable hypotheses. This aligns with the growing recognition of LLMs as tools for urban sensing and knowledge synthesis. In this role, the LLM acts as a powerful catalyst for human led scientific inquiry, offering a new method to ”operationalize” qualitative knowledge into a quantitative, empirical research program.
@@
While this study provides a foundational framework, its scope defines a clear agenda for future research. Continuous benchmarking is essential, as the quantitative performance of LLMs will evolve with new architectures [20]. More importantly, our findings underscore a critical distinction. The systematic biases and lack of causal depth suggest that current LLMs, while powerful, are not true world models. They are adept at learning the ”what” from data but struggle with the ”why”. The path toward transformative breakthroughs in urban science, particularly in areas like causal inference and policy simulation [43], will likely require a new generation of hybrid models. Future work should focus on integrating the rich, contextual knowledge of LLMs with the causal integrity of traditional urban simulation frameworks, a synergy that could unlock the next frontier of a robust and predictive urban science.
@@ 4 Method
The intelligence of large language models essentially derives from their ability to compress world knowledge acquired during training [44, 45]. To systematically evaluate the capabilities of LLMs in urban science, this study employs a framework grounded in three foundational urban theories (Fig. 5). The selection of these theories is deliberate, designed to test the LLMs’ generative fidelity across the distinct yet interconnected scales at which urban phenomena are understood: the macro-level system of cities, the meso-level internal structure of a single city, and the micro-level dynamics of a neighborhood.
@@
Substantial evidence suggests that, similar to regularities observed in natural complex systems, urban phenomena exhibit patterns and order across these multiple dimensions of size, space, and morphology [36, 39, 40, 41]. The chosen theories are canonical representations of these scaled phenomena (Table 1). Urban scaling laws capture nonlinear regularities at the macro-scale, describing how infrastructure and socioeconomic activity grow with city size. Distance decay theory reflects meso-scale spatial interaction friction and geographic reach, explaining the internal organization of cities. Finally, Jane Jacobs’ theory of urban vitality articulates how fine-grained, micro-scale morphological configurations contribute to emergent urban dynamics. Together, these theories exemplify foundational mechanisms in urban systems. They provide a comprehensive and representative testbed for evaluating the core capabilities of LLMs: from reproducing mathematically explicit laws to simulating the socio-cultural knowledge embedded in qualitative theory.
Table 1: A Multi-Scale Framework for Evaluating LLMs in Urban Science.
Scale Theory Significance in Urban Science
Macro Urban scaling law Explains universal, size-based regularities of cities as
(City system) complex systems, central to understanding urban growth
and efficiency.
Meso Urban distance decay Represents the foundational role of spatial friction and
(Internal city structure) accessibility in shaping internal urban structure and
organization.
Micro Urban vitality Provides a human-centric perspective linking the micro-design
(Block) of the built environment to social and economic outcomes;
a cornerstone of urban design.
@@ 4.1 Comparison of data distribution
First, this study directly compares the data distributions of generated and real-world samples. Since LLM-generated values without geographic prompts often differ from real samples by orders of magnitude, this study applies equal-width binning and relative frequency normalization to eliminate differences in data scale, converting the raw sample counts within each bin into frequencies. The Mean Absolute Error (MAE) of bin-level frequencies is then used to measure the average difference in relative frequency between real and generated data across bins, serving as an indicator of how well the generated data approximates the overall distribution trend of the real data. This is calculated by summing the absolute differences of mean frequencies across all bins, thereby capturing the degree of deviation in central tendency between the two distributions. For each bin, the average frequency is computed separately for the real data (R) and the generated data (G). $N$ is the number of bins, $\mu_{R,i}$ is the average relative frequency of R in the ith bin, $\mu_{G,i}$ is the average relative frequency of G in the i th bin.
@@
Secondly, based on the urban data generated by LLMs and the real data of Chinese cities, this study measures the pairing difference between the two data and three discrete distributions by Jensen-Shannon Divergence (JSD). Include ’distance between real sample and real sample data distribution’ (R-R), ’distance between real sample and generated sample data distribution’ (R-G), ’Distance between generated sample and generated sample data distribution’(G-G). Taking the calculation of average divergence of R-G as an example, there are $n$ true distributions $R_{i}$ and $m$ generated distributions $G_{j}$.
@@
$\text{JSD}(R,G)$ $=\frac{1}{2}\text{KL}(R\|\frac{1}{2}(R+G))+\frac{1}{2}\text{KL}(G\|\frac{1}{2}(R+G))$ (4)
@@
4.2 Urban scaling law
For Zipf’s Law between cities, the relationship typically follows a power-law function where $\beta\approx-1$, meaning that city size (e.g., population) and city rank follow an inverse relationship. In this equation, $\log_{10}(a)$ represents the intercept, corresponding to the constant $a$ in the power-law function, while the slope $\beta$ determines the exponent of the power-law relationship. Here, $y$ represents population size, and $x$ denotes city rank. The model incorporates several statistical parameters: $n$ refers to the number of independent trials fitting the theoretical equation, $g$ represents the number of generated data samples per trial, and $m$ denotes the parameters of the theoretical equation (such as $a$ and $\beta$ in Zipf’s Law between cities). The standard deviation of $m$ across $n$ trials reflects variability in parameter estimation, while the mean value of $m$ provides an overall estimate. A narrower confidence interval (CI) for $\beta$ suggests a more precise estimate. Additionally, the coefficient of variation (CV) for Zipf’s Law parameters is computed as the geometric mean of the CV for $a$ and $\beta$, assuming that these parameters are independent and without considering their correlations.
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (7)
The urban scaling law in cities describes how the infrastructure and socio-economic factors in a city change with the growth of the city scale. It is the scaling law of urban infrastructure and urban social factors derived by Bettencourt in 2013 based on the Human Interaction model. $Y$ represents the infrastructure or socio-economic elements in the city. $\text{CV}_{\text{Urban}}$ represents the coefficient of variation of the parameters of the urban scaling law in cities under $n$ independent experiments.
$\log_{10}(y)=\log_{10}a+\beta\log_{10}(\text{Population})$ (8)
The understanding of Zipf’s Law and urban scaling law by the LLMs was tested for inter-city and intra-city models, respectively. By inputting the appropriate prompt, the large model is used to generate city samples, including city serial number, population, road length, and GDP, and the theoretical equations of inter-city model and urban internal model are fitted respectively, and the goodness of fit and regression coefficient are observed, and the differences between different large models are compared. And compare it with the value of empirical research [38]. Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
Table 2: Prompt and specific description for urban scaling law.
Prompt Description
Urban scaling law Generate the dataset containing 100 cities in a country. Format the output as: CityName1, Population1, Infrastructure volume1, GDP1 CityName2, Population2, Infrastructure volume2, GDP2 Output exactly 100 lines without any additional text or explanations. - City Name - Population - Infrastructure volume: Total road miles - GDP: all Gross Domestic Product of the city in one year
4.3 Urban distance decay
The term distance decay is used to describe the diminishing intensity of interaction between two locations as the distance between them increases. It refers to the phenomenon where the strength of connections — especially in economic activities — tends to weaken with greater spatial separation [47]. At its core, the distance decay law posits that interactions between geographic entities are inversely related to distance; under otherwise equal conditions, such interactions typically follow an inverse-square relationship with distance. As a concept that reflects the decline of influence over space, distance decay is closely related to Tobler’s First Law of Geography [48]: ’Everything is related to everything else, but near things are more related than distant things’.
In urban geography, cities and towns often have clearly defined centers, and most urban attributes — such as population density, land value, and the quality or quantity of services — tend to decrease with increasing distance from the center [47]. Due to spatial agglomeration, urban spatial structures frequently exhibit a pronounced ’center-periphery’ pattern, where the intensity of urban elements decays with increasing distance from the city core. The most typical manifestation of this is the spatial decay of population density. Common quantitative models include the negative exponential model [49] and the inverse power model [50]. For urban land use density, the inverse S-function model has also been proposed to characterize its spatial decay [51].
For urban land density, the inverse S-function model describes a decaying distribution pattern that follows an inverse S-shaped curve as one moves outward from the city center. In this model, $L$ represents the saturation density parameter, indicating the maximum possible land density (i.e., the asymptote of the curve). $k$ denotes the growth rate parameter (1/km), which reflects how rapidly the density transitions from low values to the saturation level; the larger the value of $k$, the steeper and more abrupt the transition. $x_{0}$ is the center point parameter (km), representing the location at which the land density reaches half of its maximum value — essentially, the inflection points of the curve. $\mathrm{CV}_{\mathrm{DLand}}$ represents the coefficient of variation of the related parameters of land density attenuation in the distance decay theory under $n$ independent experiments.
$f(r)=\frac{1-c}{1+e^{a\left(\frac{2r}{D}-1\right)}}+c$ (9)
The extent to which current LLMs understand the spatial distribution of population density and land density according to the distance decay law was assessed. By providing appropriate prompts, LLMs simulated a mirage city with 100 concentric zones and generated samples that included zone identifiers, population density, and land density attributes for each zone. Then the study fitted theoretical equations for both population and land density to the generated data, evaluating the goodness of fit and regression coefficients, and compared these values with those found in empirical studies [52]. Additionally, the study examined the differences across various LLMs. Each model conducted 100 independent trials to generate data samples, and finally reported the averaged results across these repetitions.
Table 3: Prompt and specific description for urban distance decay theory.
Prompt Description
4.4 Jane Jacobs’ urban vitality
Urban vitality theory is a classic urban theory put forward by Jacobs [36]. She pointed out that urban vitality often depends on four key elements: concentration (population density), functional diversity (architectural function mixture), contact opportunity (short block), and need for aged buildings (proportion of old buildings). These factors together constitute the foundation of urban vitality, among which population density and architectural function mixing are the core driving forces, while short blocks and old buildings enhance urban vitality by increasing contact opportunities and reducing operating costs. However, in recent years, there have been disputes about Jacobs’ views on high-rise buildings, and some studies believe that it has a positive impact on urban vitality [53].
Although Jacobs’ theory is widely accepted, empirical tests remain limited and inconsistent. The reasons for uncertainty are the robustness of the research design and the interpretability of the theory to cross-geographical areas. According to Jane Jacobs’ urban vitality Theory, there is a significant positive correlation between urban vitality and given factors. $P$ stands for Population Density, $B$ stands for Building Mix Index, $S$ stands for Short Block, $A$ stands for Aged Building, and $T$ stands for Tall Building.
$P=f(P,B,S,A,T,e)$ (10)
The understanding of Jane Jacobs’ urban vitality theory by the LLMs was tested. By inputting the appropriate prompt, the large model is used to generate urban samples, including land serial number, population density, building mixing degree, street intersection density, the proportion of old buildings and building height [53], and LLM is called as a resident for the second time to judge the livability of a block, so as to be the representative value of vitality, fit the theoretical equation, observe the goodness of fit and regression coefficient, and compare the differences of different large models. And compared with the numerical value of empirical research [54]. Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
Table 4: Prompt and specific description for urban vitality theory.
Prompt Description
Urban vitality theory – generation Generate attributes for 100 blocks in a city. Format the output as: Block Name1, Population Density1, Building Mix Index1, Short Block1, Aged Building1, Tall Building1 Block Name2, Population Density2, Building Mix Index2, Short Block2, Aged Building2, Tall Building2 Output exactly 100 lines without any additional text or explanations. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. Return value 0 and 1. - Short Block: Total number of street intersections divided by the area of the block. - Aged Building: The rate of old buildings on each block. Return value 0 and 1. - Tall Building: The average building heights on each block.
Urban vitality theory – evaluation You are a resident of a city. Assign a livability score between 0 and 1 for each block, 0 represents the least livable and 1 represents the most livable. Evaluate each block based on the following attributes: - Block name. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. - Short Block: Total number of street intersections divided by the area of the block. - Aged Building: The rate of old buildings on each block. - Tall Building: The average building heights on each block.

### v3

@@ 3 Discussion
This study reveals a fundamental duality in the capabilities of current GenAI models within urban science. They excel at reproducing the abstract, mathematical forms of established statistical theories but falter in capturing the complexity of real-world urban systems. This pattern of oversimplification extends directly to the visual domain. The systematic evaluation across different scales demonstrates that while GenAI models can generate data that is structurally ”form-similar” to urban theories, this apparent competence masks significant deficiencies. In the statistical realm, this manifests as a lack of diversity and systematically distorted parameters (e.g., underestimated scaling exponents). In the visual realm, it appears as a bias against perceptual ambiguity (agreeing with human ’equal’ judgments in only 2.1% of cases) and a failure to capture the full variance of global urban environments.
@@
The initial finding that GenAI models consistently generate ”mirage cities” with low diversity and systematically distorted parameters (e.g., underestimated scaling exponents) points to a core limitation. This phenomenon, where models produce archetypal representations of cities, a tendency sometimes described as ”group flattening” [38], aligns with observations of geographical and social bias in other domains [31, 32]. This interpretation suggests that these distortions are symptomatic of the GenAI’s nature as statistical pattern matchers rather than causal world models. Lacking a deep, generative understanding of the mechanisms that produce urban phenomena, they instead reproduce the most dominant statistical patterns from their training data, smoothing over the granular heterogeneity and outliers that are often crucial drivers of real world dynamics. This tendency, also noted in other applications of GenAI models for generating synthetic populations [30], poses a significant risk for empirical research that relies on capturing the full spectrum of urban complexity.
@@
The investigation into data fidelity reveals that GenAI models are malleable tools sensitive to user interaction and model specific characteristics. Our finding that prompt engineering can partially mitigate deviations by steering the model toward a specific context underscores the importance of the human AI interface in scientific applications [29, 30, 39]. The observed performance differences across the five evaluated GenAI models suggests that future improvements in data fidelity will depend on both more sophisticated interaction protocols and the rapid advancements in the models themselves [20]. Beyond these interactive refinements, our successful implementation of a post-hoc calibration procedure using Optimal Transport offers a crucial insight: the raw output of an GenAI model should be seen as a structurally rich yet uncalibrated starting point. This hybrid approach effectively overcomes the limitations of relying solely on the model’s internal world representation, demonstrating a practical path toward creating high-fidelity synthetic datasets.
@@
A distinct capability of GenAI models is revealed when they are tasked with navigating less formalized, qualitative theories like urban vitality. Here, the GenAI is prompted to simulate a ”social consensus” by synthesizing the vast amount of information about cities embedded in its training data, from academic texts to public discourse. The result is the generation of a plausible, structured set of relationships that can provide urban scientists with novel, testable hypotheses. This aligns with the growing recognition of GenAI models as tools for urban sensing and knowledge synthesis. In this role, the GenAI acts as a powerful catalyst for human led scientific inquiry, offering a new method to ”operationalize” qualitative knowledge into a quantitative, empirical research program.
The GenAI models also unveil a novel scientific opportunity. Their tendency to generate data by guidance of prompts with controlling variables in experiments, makes them powerful instruments for theoretical exploration [28, 29, 30, 31, 32]. Urban scientists can now generate vast, fine grained datasets under specific theoretical assumptions, creating virtual laboratories to explore a theory’s logical consequences or test its sensitivity to certain parameters. This moves beyond simply validating models against existing data and toward a new mode of inquiry focused on understanding the internal consistency and emergent properties of theories themselves. This approach directly addresses the challenge of how AI can generate new knowledge for urban studies, by providing a scalable method to rigorously interrogate theoretical constructs in a controlled environment. However, while our structured prompting strategy enhanced the diversity of the generated outputs, a fundamental limitation persists: the current methodology is incapable of synthesizing a truly de novo virtual city and is instead constrained to generating data corresponding to random, extant locations on Earth.
While GenAI models can be prompted to generate data for hypothetical or counterfactual scenarios, interpreting these outputs as scientifically valid predictions is challenging. For example, correcting known biases in generated data to test unknown scenarios where data cannot be obtained in the real world. The primary obstacle is the ”black box” nature of the models. We cannot audit their internal reasoning process. Furthermore, the systematic biases we measure in one context, such as under baseline conditions, are not guaranteed to be stable when the model is prompted with a novel, out-of-distribution scenario. Therefore, these counterfactual generations are most valuable when used to produce structured and testable hypotheses that guide subsequent empirical research.
@@
While this study provides a foundational framework, its scope defines a clear agenda for future research. Continuous benchmarking is essential, as the quantitative performance of GenAI models will evolve with new architectures [20]. More importantly, our findings underscore a critical distinction. The systematic biases and lack of causal depth suggest that current GenAI models, while powerful, are not true world models. They are adept at learning the ”what” from data but struggle with the ”why”. The path toward transformative breakthroughs in urban science, particularly in areas like causal inference and policy simulation [40], will likely require a new generation of hybrid models. These findings position GenAI models as powerful engines for synthesizing urban scientific data and exploring the socio-cultural patterns for scientific inquiry. Future work should focus on integrating the rich, contextual knowledge of GenAI models with the causal integrity of traditional urban simulation frameworks, a synergy that could enable new avenues for a robust and predictive urban science.
@@ 4 Method
The intelligence of GenAI models essentially derives from their ability to compress world knowledge acquired during training [41, 42]. To systematically evaluate the capabilities of GenAI models in urban science, this study employs a framework that tests their generative fidelity against two core domains of urban theory: (1) statistical laws based on symbolic data (different dimensions with scale, space, and morphology), and (2) cognitive laws based on perceptual data (such as street-view images in urban scenes).
Figure 6: AI4US, the proposed framework utilizing GenAI models to advance urban science.
4.1 Urban theories based on generating symbolic data
@@
For the first domain, we selected three foundational theories designed to test the GenAI’s abilities across the distinct yet interconnected scales at which urban phenomena are understood: the selection of these theories is deliberate, designed to test the GenAI’s generative fidelity across the distinct yet interconnected scales at which urban phenomena are understood: the macro-level system of cities, the meso-level internal structure of a single city, and the micro-level dynamics of a neighborhood (Supplementary Note 2). Substantial evidence suggests that, similar to regularities observed in natural complex systems, urban phenomena exhibit patterns and order across these multiple dimensions of size, space, and morphology [36, 43, 44, 45]. The chosen theories are canonical representations of these scaled phenomena (Table S1). Urban scaling laws capture nonlinear regularities at the macro-scale, describing how infrastructure and socioeconomic activity grow with city size. Distance decay theory reflects meso-scale spatial interaction friction and geographic reach, explaining the internal organization of cities. Finally, Jane Jacobs’ theory of urban vitality articulates how fine-grained, micro-scale morphological configurations contribute to emergent urban dynamics. Together, these theories exemplify foundational mechanisms in urban systems. They provide a comprehensive and representative testbed for evaluating the core capabilities of GenAI models: from reproducing mathematically explicit laws to simulating the socio-cultural knowledge embedded in qualitative theory.
4.1.1 Urban Scaling Law
As a foundational concept in complex systems, urban scaling laws describe the power-law relationship between a city’s size and its various metrics, serving as functional models for understanding cities as complex systems. These laws are broadly categorized into inter-urban processes (interactions between cities) [46], such as Zipf’s Law, and intra-urban processes (dynamics within a city) that explore how human interactions and spatial geometry contribute to scaling. Zipf’s Law posits an inverse relationship between a city’s population and its rank.
The intra-urban scaling law relates socioeconomic or infrastructural metrics $y$ to city population.
$\log_{10}(y)=\log_{10}a+\beta\log_{10}(\text{Population})$ (2)
$\log_{10}(a)$ represents the intercept, while the slope $\beta$ is the critical scaling exponent. Empirical evidence [38] suggests two distinct patterns for intra-urban scaling: socioeconomic activities (e.g., GDP) exhibit super-linear scaling ($\beta>1$), implying increasing returns to scale, while infrastructure metrics show sub-linear scaling ($\beta<1$), indicating greater efficiency in larger cities. To test these laws, GenAI models were prompted to generate a dataset of 100 cities with attributes for population, infrastructure, and GDP.
4.1.2 Urban Distance Decay
This theory addresses the meso-scale structure of cities, positing that the intensity of interaction between two locations diminishes with increasing distance, a concept closely related to Tobler’s First Law of Geography [47, 48]. This typically results in a pronounced ”center-periphery” pattern in urban spatial structures, where attributes like population density and land value decrease from the city core. While several quantitative models exist, our study focused on the inverse S-function model for characterizing the spatial decay of urban land density $f(r)$ as a function of distance ($r$) from the city center [49, 50, 51].
$f(r)=\frac{1-c}{1+e^{a\left(\frac{2r}{D}-1\right)}}+c$ (3)
The parameters define the curve’s shape: $c$ represents the minimum land density in the urban periphery (the lower asymptote); $a$ is the growth rate, controlling the steepness of the density decay; and $D$ is a scaling parameter related to the city’s spatial extent, where the inflection point of the curve occurs at the distance $r=D/2$. To assess the models’ grasp of this concept, GenAI models were prompted to generate land density values for 100 concentric urban zones.
4.1.3 Jane Jacobs’ Urban Vitality
At the micro-scale, we tested the GenAI’s ability to operationalize Jane Jacobs’ classic theory of urban vitality [35]. The theory posits that vitality depends on four key elements: concentration, functional diversity, contact opportunity (short blocks), and aged buildings, where population density and building mix are considered core driving forces [52]. Despite its wide acceptance, empirical tests of the theory remain limited and inconsistent [huang2024b], and debates persist around certain elements, such as the role of tall buildings.
$P=f(P,B,S,A,T,e)$ (4)
Here, vitality is a function of Population Density ($P$), Building Mix Index ($B$), Short Block ($S$), Aged Building ($A$), and Tall Building ($T$). We employed a two-step process: first, GenAI models were prompted to generate data for 100 urban blocks across these five key attributes; second, the models were prompted to act as a ”resident” and assign a ”livability score” to each block as a proxy for vitality, allowing us to test whether the models’ inferred relationships were consistent with Jacobs’ core tenets.
4.2 Urban perception through generating perceptual data
For the second domain, we evaluated the GenAI’s multimodal capabilities in urban visual perception. We focused on human-scale street view imagery (SVI), a mature benchmark representing the daily urban life. The evaluation used a ”Generate-Evaluate-Calibrate” paradigm grounded in the six perceptual indicators from the MIT Place Pulse project [37, 53, 54, 55, 56, 57] (Table S2). Our methodology involved a three-phase process to validate the AI as a visual data generator, a perceptual judge, and a tool for causal inference.
Figure 7: A computational framework for advancing urban perception through perceptual data. Edit text and images through Nano Banana.
First, to validate the AI’s data generation capability, the generated human-scale streetscapes were assessed for realism and diversity. Realism was evaluated by human experts, who rated 100 images from AI generation and 100 images from Place Pulse on a 1-10 scale (Supplementary Experiment 9). The model’s fidelity was then quantified by comparing the score distributions against real photographs using an independent samples t-test. Dataset diversity was ensured via a structured-prompting strategy. The effectiveness of this strategy was then quantitatively validated by using CLIP embeddings to compare the average semantic distance $D_{CLIP}$ across three distinct image sets: a real-world street view benchmark from Place Pulse, a baseline set generated from simple prompts, and our proposed structured-prompt set.
$D_{CLIP}=\frac{2}{N(N-1)}\sum_{i=1}^{N}\sum_{j=i+1}^{N}\sqrt{1-\frac{e_{i}\cdot e_{j}}{\|e_{i}\|\|e_{j}\|}}$ (5)
Second, to validate the AI’s perceptual judgment, its pairwise choices on original Place Pulse images were compared against the source dataset’s human choices using Kappa.
$K=\frac{P_{0}-P_{e}}{1-P_{e}}$ (6)
Furthermore, to assess if the AI could identify influential visual factors in alignment with human patterns [57], we fit a multiple linear regression model correlating the AI’s perception scores ($Y_{p}$) with the pixel proportions of 150 visual elements ($X_{k}$) from semantic segmentation.
$Y_{p}=\beta_{p,0}+\sum_{k=1}^{K}\beta_{p,k}X_{k}+\varepsilon_{p}$ (7)
Third, to validate the AI’s utility for causal inference, we performed thematic interventions on baseline images to create ”before-and-after” pairs. These interventions were designed around three interpretable categories: Natural, Traffic, and Built Elements. The causal effect of each category was then quantified by calculating the average change in perception scores ($\Delta S_{p,c}$) between the variant and baseline images for each perception type.
$\Delta S_{p,c}=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}(S_{p,i}^{\text{variant}}-S_{p,i}^{\text{baseline}})$ (8)
Where $\Delta S_{p,c}$ is the average change in score for perception $p$ under category $C$, $S_{p,i}^{\text{variant}}$ and $S_{p,i}^{\text{baseline}}$ are the scores for a single image pair $i$, and $N_{c}$ is the total number of image pairs in that category.
@@ 4.3 Comparison of data distribution
First, this study directly compares the data distributions of generated and real-world samples. Since generated values without geographic prompts often differ from real samples by orders of magnitude, this study applies equal-width binning and relative frequency normalization to eliminate differences in data scale, converting the raw sample counts within each bin into frequencies. The Mean Absolute Error (MAE) of bin-level frequencies is then used to measure the average difference in relative frequency between real and generated data across bins, serving as an indicator of how well the generated data approximates the overall distribution trend of the real data. This is calculated by summing the absolute differences of mean frequencies across all bins, thereby capturing the degree of deviation in central tendency between the two distributions. For each bin, the average frequency is computed separately for the real data (R) and the generated data (G). $N$ is the number of bins, $\mu_{R,i}$ is the average relative frequency of R in the ith bin, $\mu_{G,i}$ is the average relative frequency of G in the i th bin.
@@
Secondly, based on the urban data generated by GenAI models and the real data of Chinese cities, this study measures the pairing difference between the two data and three discrete distributions by Jensen-Shannon Divergence (JSD). Include ’distance between real sample and real sample data distribution’ (R-R), ’distance between real sample and generated sample data distribution’ (R-G), ’Distance between generated sample and generated sample data distribution’(G-G). Taking the calculation of average divergence of R-G as an example, there are $n$ true distributions $R_{i}$ and $m$ generated distributions $G_{j}$.
@@
$\text{JSD}(R,G)$ $=\frac{1}{2}\text{KL}(R\|M)+\frac{1}{2}\text{KL}(G\|M)$ (12)
$\text{where }M$ $=\frac{1}{2}(R+G)$
@@
4.4 Post-hoc calibration procedure for data distribution
To address the systematic bias and lack of diversity observed in raw generated data, we introduce a post-hoc calibration procedure (detailed explanations refer to Supplementary Note 3). This challenge of aligning a generated (’source’) distribution with an empirical (’target’) distribution is a fundamental problem in machine learning, analogous to domain adaptation [58, 59] or batch effect correction [60]. While prior work in generated data synthesis has focused on prompt engineering or simple rule-based adjustments [26, 30, 31], these methods may not correct for systemic distributional shifts. We therefore employ an approach grounded in Optimal Transport (OT) theory [58, 61] for its ability to transform an entire dataset while preserving its core structure. The procedure learns an optimal transport map $T$ that transports samples from the preprocessed generated distribution ($P_{G}$) to the preprocessed real data distribution ($P_{real}$). This map is then applied to the full generated dataset ($x_{i}$) to produce a calibrated set ($D_{\text{calib, scaled}}$):
$D_{\text{calib, scaled}}=\{T(x_{i})|x_{i}\in D_{G,\text{scaled}}\}$ (14)
Here, $x_{i}$ is the set of generated data points, and T is the learned mapping function. The distance between distributions is quantified using the Wasserstein distance ($W_{1}$), which represents the minimum ”cost” required to transform one distribution into another:
$W_{1}(P_{real},P_{G})=\inf_{\gamma\in\Pi(P_{real},P_{G})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|d\gamma(x,y)$ (15)
Where $\Pi(P_{real},P_{G})$ is the set of all possible joint distributions of the source $P_{G}$ and target ($P_{real}$) distributions. The efficacy of this OT approach was validated against several baselines, including raw GenAI output, random noise, and data synthesized via KDE. This comprehensive correction is achieved because the OT procedure learns a mapping that transforms the entire joint distribution of the multi-dimensional data. By adjusting the data structure holistically, rather than treating each variable independently, it inherently corrects the inter-variable correlations that define the scaling exponents.